OpenStudy (anonymous):

5 2/5 write the fraction as a percent.

OpenStudy (brinazarski):

Okay, so first- do you know how to turn a regular fraction into a percent?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

\[5\frac{ 2 }{ 5 }=\frac{ 5*5+2 }{ 5 }=\frac{ 25+2 }{ 5 }=\frac{ 27 }{ 5 }\]

OpenStudy (brinazarski):

You divide the numerator by the denominator, and then you multiply that by a 100.

OpenStudy (brinazarski):

Do you understand what itsmylife just did?

OpenStudy (anonymous):

Yes. I just did that..it came to 40. but its not a choice

OpenStudy (brinazarski):

The answer is not 40... what do you get when you divide 27 by 5?

OpenStudy (anonymous):

5.4

OpenStudy (brinazarski):

Now the fraction is in decimal form, so multiply that by a 100 and you'll have your percentage!

OpenStudy (anonymous):

Thank you!! :D

OpenStudy (brinazarski):

You're welcome :)